OpenStudy (anonymous):

write in vertex form= 2x^2 + 6x +1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes what

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im deporting u

OpenStudy (anonymous):

deporting me?

OpenStudy (anonymous):

back to china? ono

OpenStudy (ranga):

2x^2 + 6x +1 = 2(x^2 + 3x + 1/2). Complete the square of x^2 + 3x + 1/2

OpenStudy (anonymous):

thank you

OpenStudy (ranga):

You are welcome.

OpenStudy (anonymous):

how do you get the answer for completing the square? like you do 3 over 2, then how do you square that again

OpenStudy (ranga):

2x^2 + 6x +1 = 2(x^2 + 3x + 1/2) = 2{ (x+3/2)^2 - 9/4 + 1/2 } = 2{ (x+3/2)^2 - 9/4 + 2/4 } = 2{ (x+3/2)^2 - 7/4 } = 2(x+3/2)^2 - 7/2 is the vertex form. The vertex is (-3/2, -7/2).